### v1

# Human-in-the-Loop Optimization with Model-Informed Priors

Abstract.
Human-in-the-loop optimization identifies optimal interface designs by iteratively observing user performance. However, it often requires numerous iterations due to the lack of prior information. While recent approaches have accelerated this process by leveraging previous optimization data, collecting user data remains costly and often impractical. We present a conceptual framework, Human-in-the-Loop Optimization with Model-Informed Priors (HOMI), which augments human-in-the-loop optimization with a training phase where the optimizer learns adaptation strategies from diverse, synthetic user data generated with predictive models before deployment. To realize HOMI, we introduce Neural Acquisition Function+ (NAF+), a Bayesian optimization method featuring a neural acquisition function trained with reinforcement learning. NAF+ learns optimization strategies from large-scale synthetic data, improving efficiency in real-time optimization with users. We evaluate HOMI and NAF+ with mid-air keyboard optimization, a representative VR input task. Our work presents a new approach for more efficient interface adaptation by bridging in situ and in silico optimization processes.
Human-in-the-loop optimization, user modeling, user inference, Bayesian optimization, meta-learning.
$c$s: Human-centered computing Interaction techniquesccs: Computing methodologies Machine learning
Figure 1. The general steps of our HOMI framework. Step 1: Model selection Designers or developers select a model, parameterized by $\theta$, that are relevant to the target task, and optionally fit the model’s parameters (e.g., Fitts’ law, with parameters $a$ and $b$) using minimal user data to better reflect the target context. Step 2: Synthetic user generation — A diverse set of synthetic users is created by sampling different parameter settings $\theta_{1},\theta_{2},\theta_{3},\dots$} from the fitted model’s parameter distribution. Step 3: Meta-BO training — The meta-optimizer interacts extensively with these synthetic users to learn efficient strategies for optimizing across user variability. Step 4: Deployment — The trained meta-BO is deployed with real users and quickly adapts to individual performance or preferences using the learned prior experience.
1. Introduction
Efficient and effective interface optimization for different users and contexts is a long-standing goal in human-computer interaction (HCI) (Norcio and Stanley, 1989; Zhai et al., 2002; Rao, 2019; Lavie and Meyer, 2010). Model-based optimization approaches use computational optimizers to search for interface designs that maximize predicted user performance, based on established models. The challenge of this approach is that practical models are scarce, and they are not able to capture specific users or contexts without model parameter inference (Beaumont et al., 2002; Moon et al., 2023), a known challenging step; therefore, this approach can not scale up easily or achieve true personalization. In recent years, human-in-the-loop optimization (HILO) emerged for its higher adaptability and flexibility (Lee et al., 2020; Khajah et al., 2016; Chiu et al., 2020; Kim et al., 2017). Unlike model-based methods, HILO operates directly on actual users: the optimizer proposes a design candidate, users interact with it, their performance is evaluated, and the optimizer updates the design candidate based on a series of observations. Various computational techniques, such as Bayesian optimization (Koyama et al., 2017; Liao et al., 2023; Brochu et al., 2010; Dudley et al., 2019), evolutionary and genetic algorithms (Battiti and Passerini, 2010; Meng and Wu, 2008), and heuristic-based techniques (Lee et al., 2020) have been explored in human-in-the-loop optimization. Among these, Bayesian optimization (BO) stands out due to its better performance and minimal assumptions about the task (Frazier, 2018). HILO with BO has been successfully applied in domains such as novel interaction techniques (Liao et al., 2020), design tools (Khajah et al., 2016; Brochu et al., 2010), and wearable interfaces (Liao et al., 2024a; Kim et al., 2017). Despite its adaptability, HILO often suffers from slow convergence due to a lack of prior information; it typically begins with a random search, requiring many iterations to find a good solution (Wang et al., 2022; Pu et al., 2024). This “cold-start” problem limits its deployment in real-time applications or dynamic environments.
Recent work has attempted to mitigate this limitation by incorporating prior data from past users to initialize or guide the optimization process, a concept known as transfer learning or meta-learning. Prior data was used for example to construct a prior surrogate model (Liao et al., 2024b), to construct a series of prior models and aggregate them via a weighted-sum approach (Liao et al., 2024a), or, most recently to construct an adaptive surrogate model using Bayesian neural networks (Liao et al., 2025) on which the HILO approach was then trained before deployment to end-users. These works demonstrated that augmenting HILO with prior knowledge is a viable path toward solving the cold-start issue. However, a key bottleneck remains: the reliance on real human data. Prior transfer- or meta-learning–based methods require data from at least 10 users to be effective (Liao et al., 2024b, a), which can be prohibitively expensive or impractical. Moreover, any changes to the interaction task or interface require fresh rounds of data collection, severely limiting scalability. This motivates us to ask: Is it possible to train computational optimizers for higher online efficiency, yet, without relying on real human data?
In this work, we propose a new HILO framework called Human-in-the-Loop Optimization with Model-Informed Priors (HOMI). Rather than relying on real user data for training the optimizer, HOMI leverages data from synthetic users simulated from predictive user models. This training phase allows Bayesian optimization to develop a broad understanding of the design space and learn effective adaptation strategies for diverse user contexts before the online-optimization process starts. A key advantage of using model-informed priors is that the optimizer can be trained with a wide variety of synthetic users, enabling generalization across different usage scenarios. In contrast, training with real users limits the optimizer’s knowledge to a small number of specific individuals, reducing its adaptability. At deployment, the trained optimizer switches to a human-in-the-loop mode, integrating prior knowledge with live user feedback to efficiently optimize the interface design in real time.
To realize the HOMI framework, we introduce a novel Bayesian optimization method enhanced by meta-learning; this type of approaches is commonly referred to as meta-BO in the machine learning field (Li et al., 2021; Wang, 2021). Our method, called Neural Acquisition Function+ (NAF+), builds on the work of Volpp et al. (2019), who proposed replacing acquisition functions in Bayesian optimization with neural networks trained via reinforcement learning (Kaelbling et al., 1996; Sutton and Barto, 2018). Leveraging the scalability of neural networks, NAF+ can be trained on large-scale synthetic user datasets, and incorporates two novel features tailored specifically for human-in-the-loop scenarios. First, NAF+ provides designers with enhanced flexibility in setting design objectives by dynamically aggregating multiple design objectives into a single weighted-sum objective for which individual weights can be flexibly tuned in practical applications. Secondly, NAF+ ensures robust performance across the diverse behaviors of end-users, even if their behavior varies significantly from that of synthetic users. Therefore, we propose a novelty detector that estimates how likely a user at run-time is different from the synthetic users, which is used to aggregate outputs from the neural and a standard acquisition function. We validate the effectiveness of our method through a series of synthetic tests, showing that NAF+ converges more efficiently than established meta-BO approaches, adapts its optimization strategy to varying objective weightings, and remains robust when encountering novel synthetic users.
We further demonstrate the effectiveness of our HOMI by applying NAF+ to real-time adaptation of a mid-air keyboard for direct touch input, which is an essential VR text entry technique, in a study with real users. The training involved synthetic users constructed via established models in HCI, including Fitts’ law (MacKenzie, 1992) and typing error model (Bi and Zhai, 2016). The results showed HOMI enabled by NAF+ outperforms the established baselines, including both manual keyboard setting and standard BO in that it significantly reduced the number of iterations required to personalize the keyboard.
To summarize, we make the following contributions:
  • Concept and Framework: We propose HOMI, a novel framework that trains HILO using synthetic user data generated from models, enabling scalable and efficient interface adaptation. Critically, HOMI repositions the role of user models from being optimization targets to training resources.
  • Method: We introduce NAF+, a Bayesian optimization method with a neural acquisition function trained on synthetic data, supporting dynamic objective weighting and robust generalization to unseen user behaviors.
  • Demonstration: We evaluate HOMI and NAF+ on mid-air keyboard adaptation, showing that our approach significantly outperforms standard BO and manual baselines in online adaptation scenarios.
2. Background & Related work
We review the key topics that are relevant to this work, especially focusing on (1) traditional optimization in HCI, (2) human-in-the-loop optimization (HILO), and the (3) techniques that augment Bayesian optimization with prior data for efficient interface adaptation.
2.1. Model-based Optimization in HCI
Optimization is an essential process in HCI and design, even if they are not explicitly articulated as such. In traditional user-centered design workflows, optimization is performed manually: designers generate a set of candidate designs, which are then evaluated through user testing and iteration (Abras et al., 2004). The field of computational interaction (Oulasvirta et al., 2018) aimed to integrate computational optimization methods in the design process to automatically identify optimal design candidates. This idea is not new: engineering fields have long used optimization workflows in which a solver proposes candidate designs, and these are evaluated through domain-specific models (e.g., physical, mechanical, chemical) to identify optimal solutions (Martins and Ning, 2021). This approach has proven successful in fields such as mechanical engineering (Rao et al., 2011), chemical engineering (Bhaskar et al., 2000), and architectural design (Michalek et al., 2002). HCI researchers have adopted similar methods: deploying computational solvers over user models to identify optimal interaction designs. Problems are often framed as assignment problems and solved using integer programming. Examples include menu optimization (Bailly et al., 2013; Todi et al., 2021), keyboard layout design (Feit et al., 2021; Karrenbauer and Oulasvirta, 2014), and UI layout adaptation in AR/VR environments (Lindlbauer et al., 2019; Evangelista Belo et al., 2022), and more. While this model-based approach has shown success in generating promising designs, it suffers from two key limitations: First, the application scope is highly limited by the availability of user models. Yet, practical models are scarce in real-world applications. Secondly, and more importantly, the user models represent an abstract “average user” but not any specific user under a specific context. Once deployed, such systems do not update or adapt their underlying model based on live observations, limiting their ability to adapt designs to individual needs or preferences. This work builds upon the foundation of model-based optimization but extends it into the domain of human-in-the-loop adaptation. In the HOMI framework, user models are not treated as fixed sources for optimization, but rather as generators of synthetic user data to train an optimizer. This repositions models from being optimization targets to being training resources, enabling scalable pre-training optimizers for HILO.
2.2. Human-in-the-Loop Bayesian Optimization
Human-in-the-loop optimization (HILO) is an emerging optimization paradigm in HCI that does not rely on predefined user models. Instead, HILO learns and optimizes directly from online observations, which may include user performance, preferences, or other forms of interaction feedback. The objective function $f$ is usually a user-centered performance metric (e.g., completion time, preferences, etc) that we want to optimize for (Liao et al., 2020, 2024a; Kim et al., 2017; Yamamoto et al., 2022). Each unique design candidate $x$ within the designated design space $\mathcal{X}$ results in a user-related objective function value, $f(x)$. The goal of HiLO is to identify the global optimal design $x^{*}$ that maximizes (or minimizes) the objective value $f(x^{*})$. Such optimization problem is challenging since the only way to access the user-relevant objective function is through costly user evaluations. Bayesian optimization (BO) is a popular computational method for human-in-the-loop optimization due to its minimum requirements and assumptions of the problem and reliable performances (Mockus, 1975; Shahriari et al., 2016). Within Bayesian optimization, a surrogate model, typically a Gaussian Process regression (GP), captures the properties of the target function. This model is continually refined with online function observations (the result of user evaluation), enhancing its prediction accuracy. An acquisition function utilizes the information of the surrogate model to computes the acquisition value (i.e., the worth or potential value) of a new candidate design. The candidate with the highest acquisition value is then selected for user evaluation in the subsequent iteration.
Bayesian optimization has found many human-involved applications where the sample efficiency is particularly important, such as exoskeleton systems (Kim et al., 2017, 2019; Tucker et al., 2020), wearable interfaces (Liao et al., 2023; Guo et al., 2019), haptic displays (Liao et al., 2023; Catkin and Patoglu, 2023), input techniques (Liao et al., 2024b), and design tools (Brochu et al., 2010; Koyama et al., 2017; Mo et al., 2024). Despite its robust performance and relatively better efficiency over other optimization approaches (simulated annealing, evolutionary algorithms, etc.), a key limitation of BO lies in its “cold-start” problem. Since BO generally starts without any prior knowledge of the objective function, it must rely on random exploration in the early stages, which can slow down convergence. To address this, recent work has explored incorporating transfer learning and meta-learning to accelerate the optimization process, which we will review in the next subsection.
2.3. Transfer- and Meta-learning for BO
Transfer learning (Zhuang et al., 2020) and meta-learning (Wang, 2021) are similar concepts, both referring to enhancing the model training efficiency by leveraging the data gathered from previous similar tasks. Despite their differences in deep learning, these two terms are generally interchangeable in the context of Bayesian optimization. Bai et al. (2023) provides a review of recent advancements in this topic. In the following content we will refer to the concept of “leveraging prior data to enhance the efficiency of BO” as meta-BO. One common meta-BO approach is constructing a GP model to accumulate experiences across all the previous tasks (here: end-users), typically incorporating techniques such as multi-task GP, sparse GP, and other approaches  (Bardenet et al., 2013; Bonilla et al., 2007; Swersky et al., 2013; Yogatama and Mann, 2014). However, the key bottleneck in this class of methods is scalability: GP model fitting scales cubically with the number of observations, i.e., $\mathcal{O}(n^{3})$, where $n$ is the total number of data points. The second approach is called the weighted-sum method; it stores the experience from each previous task as an individual model. When deploying on the new task, these prior models jointly guide the ongoing optimization, where their importance is determined by weights. This method offers slightly better scalability (Li et al., 2022; Schilling et al., 2016; Wistuba et al., 2018), as its computation time only increases linearly with the number of prior models. Some recent works extended this approach for wrist-based input techniques (Liao et al., 2024a) and visual appearance design (Li et al., 2025). However, this approach is still not truly scalable due to the linearly increasing complexity. Notably, a key advantage of HOMI is the unlimited synthetic data generated by the user models. Following the previous approaches, regardless of a single model or a series of models, will not fully unlock the potential of unlimited dataset. To overcome this limitation, our approach follows a third direction – replacing hand-crafted components of BO with deep neural networks, enabling scalable training with large datasets. For instance, a deep kernel (Wistuba and Grabocka, 2021) or fully replacing the surrogate and acquisition function with learned models (Hsieh et al., 2021). Our NAF+ is built upon Volpp et al. (2019), which replaces the manually crafted acquisition function with a deep-neural net, called Neural Acquisition Function (NAF). Pre-trained with data of similar tasks, NAF acquired the specific strategy to pick the next design candidates to be evaluated of this type of tasks. In our context of human-in-the-loop optimization, we define each task as optimizing for a specific user (or synthetic user) under a shared interaction or application. While each user may exhibit different behavior or preferences, we assume there exist shared patterns across users that can be learned and transferred via meta-learning. This formulation is particularly well-suited to our framework, HOMI, where synthetic users (generated from established user models) enable the creation of unlimited training data, a key advantage over traditional approaches. Note that the neural acquisition function in NAF+ is trained via reinforcement learning, which we describe in more detail in the next section.
3. Human-in-the-Loop Optimization with Model-Informed Priors
Our first contribution is the introduction of a conceptual framework, which we term HOMI. We distinguish HOMI from our proposed method (detailed in the next section): While NAF+ represents one possible realization, the HOMI framework is intended to be general and extensible, enabling future work to explore alternative implementations. Below, we outline the core components of the framework.
3.1. Core Elements of HOMI:
The HOMI framework consists of three key components: meta-Bayesian optimization (meta-BO), user models, and end-users. Meta-BO refers to a computational optimizer trained across a distribution of synthetic users for the same interaction (or application). During the optimizer training phase, meta-BO interacts with synthetic users generated from selected, relevant user models. Since these interactions are simulated, the optimizer can perform an unlimited number of evaluations to learn generalized adaptation strategies. During the deployment phase, meta-BO engages with real end-users. Its goal is to quickly infer user-specific preferences or performance characteristics, and identify optimal designs with minimal interactions. User models are parameterized functions (e.g., Fitts’ law) that simulate user behavior and performance. By sampling different model parameter settings, we can generate a diverse population of synthetic users, each representing different motor abilities, interaction speeds, or usage contexts. End-users are real human participants interacting with the system during deployment. The meta-BO observes real-time signals, such as performance metrics, interaction preferences, or behavioral data—and uses them to adaptively select new design candidates.
3.2. General Steps in HOMI:
Figure 1 illustrates the four main steps in the HOMI framework. Note that the first three steps comprise the offline training phase, while only the fourth step occurs during real-time deployment.
Step 1: Model selection. HOMI framework requires synthetic users with varying performances to train the optimizer. User models are the tool for this. In HCI, many models are proposed to capture and predict user performances for different interactions and contexts (Purificato et al., 2024); by varying the parameter settings within these models, different user performance can be simulated. Thus, the first step of our framework is to identify the models relevant to the target task; that is, a model that can consider the variation in design and predict different user performances accordingly. For example, for typing and target selection, Fitt’s law is the established model for predicting the movement time. These user models often require parameter tuning to accurately predict performance in a given setting or context. This step may involve fitting model parameters to empirical data or adopting values reported in prior literature, yielding relevant parameter distributions. This is optional: if reliable parameter estimates already exist, they can be directly reused. For instance, touchscreen touch accuracy has been modeled using a 2D Gaussian distribution with well-documented parameter values (Bi and Zhai, 2013), which can serve as a ready basis for generating synthetic users. There may be cases that there is not a relevant model yet for a novel interaction technique. In such cases, introducing novel user models will be needed.
Step 2: Synthetic user generation: After acquiring reliable model parameter estimates and ranges, we can then sample possible model parameters from the parameter distributions. This leads to a diverse set of synthetic users, each simulating distinct user characteristics.
Step 3: Meta-BO training with synthetic interactions: Before deploying to the end-users, meta-BO interacts with these synthetic users to learn how to efficiently identify optimal designs across different user profiles. Since synthetic evaluations are inexpensive, compared to human evaluations, this phase allows for large-scale training. The diversity of synthetic users further ensures the optimizer does not overfit to specific traits, but rather learns transferrable strategies for personalization.
Step 4: Human-in-the-loop optimization: The trained optimizer, meta-BO, is deployed with real users, leveraging prior experience to adapt quickly and effectively to individual needs, ultimately addressing the cold-start problem.
4. Neural Acquisition Function+
To realize the concept of HOMI, we introduce a novel meta-BO method for HCI purposes: Neural Acquisition Function+ (NAF+). The central idea behind NAF+ is to replace the manually designed acquisition function in Bayesian Optimization with a neural network, trained on synthetic user data. This enables the optimizer to learn efficient adaptation strategies prior to deployment, resulting in faster convergence during real-time human-in-the-loop optimization. Our NAF+ is built upon a key previous work, Neural Acquisition Function (NAF) (Volpp et al., 2019), and extends it with two novel features that offer designers flexibility and ensure robust performance for real-world users. In this section, we first introduce the necessary preliminaries: standard acquisition functions in BO, reinforcement learning (the core training method behind NAF), and the original NAF formulation. We then present the core contributions and architectural enhancements of our proposed NAF+.
4.1. Preliminaries
To better introduce our proposed method, we briefly review three foundations: standard acquisition functions and their limitations, reinforcement learning, and the Neural Acquisition Function (NAF). Our NAF+ builds directly on these concepts and implementations.
1. Acquisition functions in BO
As reviewed in Related Work, BO has two core components: a surrogate model that approximates the target objective function (typically, GP) and an acquisition function that guides the search for the next design candidate to evaluate. The primary goal of BO is to efficiently identify the optimal point (i.e., design) with a minimal number of evaluations. The acquisition function leverages information from the surrogate model (specifically, the predicted mean and predicted variance for each candidate design) to compute an acquisition value. This value represents the potential of evaluating that candidate in the next iteration. The acquisition function is computed across the design space, and the candidate with the highest acquisition value is selected for evaluation. However, a shared limitation of the conventional acquisition functions is that they are not task-specific: their heuristics are fixed and do not adapt based on prior experience or the characteristics of a given domain. This limits their efficiency, especially in human-involved tasks where evaluation cost is high and adaptation must be fast. This motivates our transition toward learned acquisition functions, where the search strategy itself is learned from prior data–collected from other similar users or, in our case, synthetic users–tailored to the structure of the problem and capable of generalizing across users or tasks.
2. Deep reinforcement learning
The core idea of Neural Acquisition Function (NAF) is to learn a task-specific acquisition function using a neural network. Instead of relying on hand-designed heuristics like Expected Improvement or UCB, NAF aims to encode an adaptive search strategy that is tailored to a particular class of optimization tasks. A key challenge in learning such an acquisition function is the absence of ground truth labels — i.e., there is no “correct” acquisition value to supervise the training. The predictions are inherently uncertain, noisy, and task-dependent. To overcome this, NAF is trained using reinforcement learning (RL). Specifically, the neural network generates acquisition scores across the design space. The design candidate with the highest score is selected for evaluation. The outcome of this evaluation — i.e., the observed objective function value — is then converted into a reward signal: promising evaluation results yield positive rewards, while poor results generate negative feedback. Over time, the model learns to associate high acquisition values with promising candidates from current status, refining its search policy. Viewed from an RL perspective, meta-BO acts as an RL agent, and the neural acquisition function (NAF) serves as its policy model, determining which action (i.e., which design candidate) the agent should take in each iteration given the GP’s status. This formulation enables task-level learning, where the acquisition policy improves through repeated interaction with synthetic training tasks. Below, we briefly review the concept of RL and the corresponding method for training NAF.
RL is a machine learning paradigm that trains an agent to develop a policy for taking decisions to maximize the reward signals within a given environment. The environment is usually formulated as a Markov decision process. The state that the agent observed at time step $t$ is usually denoted as $s_{t}\in\mathcal{S}$, where $\mathcal{S}$ is the state space. The agent’s action at time step $t$ is denoted as $a_{t}$, which is determined by its policy $\pi:\mathcal{S}\rightarrow\mathcal{P(A)}$, where $\mathcal{P(A)}$ represents the set of probability distributions over the action space $\mathcal{A}$. In this framework, the policy is probabilistic, meaning that for any state $s\in\mathcal{S}$, $\pi(s)$ specifies a distribution over possible actions. The action taken ($a_{t}$) will let the agent move onto the next state ($s_{t+1}$). Such a transition is also mostly based on a probability distribution, denoted as $p(s_{t+1}|s_{t},a_{t})$. Accompanying the transmission from $s_{t}$ to $s_{t+1}$, the agent receives a reward signal, denoted as $r_{t}$, which is determined from a reward function ($r$). We can further denote the reward received at $t$ as such, $r_{t}=r(s_{t},a_{t},s_{t+1})$. The agent learns the policy via trials and errors, aiming to maximize the cumulative discounted future reward, $\eta(\pi)$, based on its policy ($\pi$). With the reward discount factor ($\gamma\in(0,1]$), the initial state ($s_{0}$), and the episode length $T$, we can formulate $\eta(\pi)$ as such: $\eta(\pi)=\mathbb{E}_{\pi}\Bigl[\sum_{t=1}^{T}\gamma^{t-1}r_{t}\,\Bigl|\,s_{0}\Bigr]$. Deep RL is an advancement in RL that incorporates deep neural networks, to handle high-dimensional input spaces more effectively. Among all the deep RL algorithms, NAF is trained via Proximal Policy Optimizations (PPO) (Schulman et al., 2017), a common policy gradient-based algorithm. Our NAF+ is built upon Volpp et al. (2019), which also applied PPO as the training algorithm.
3. Neural Acquisition Function (NAF)
NAF keeps GP as the surrogate model of the task, allowing it to model the properties of the current optimization based on the real-time observations. However, instead of hand-crafted heuristic-based acquisition functions, NAF employs a neural network to generate the acquisition values (as introduced earlier). In the original paper, NAF (the model itself) at timestep $t$ as $\alpha_{t,\theta}$, where $\theta$ indicates that it is parameterized by a neural network. Similar to standard acquisition functions, NAF takes the information of the GP model as the input. However, a GP is a continuous function, but NAF is a deep neural network, whose input has a fixed dimension. Following Volpp et al. (2019), the design space $\mathcal{X}$ is discretized into a set of points $x_{s}$}. For each candidate $x$, NAF receives the GP’s predicted mean $\mu_{t}(x)$ and standard deviation $\sigma_{t}(x)$ as input features. Since the GP evolves with incoming observations, these features are dynamically updated at every iteration. In the original implementation, the set of $x_{s}$ was selected dynamically at each step. Note, to simplify training and ensure stable input representation, in our NAF+, we adopt a fixed set of $x_{s}$ throughout the optimization process. Additionally, to help the model account for optimization progress and urgency, NAF includes a budget-aware input: the percentage of optimization completed, calculated as the current iteration index divided by the total allowed iterations. This allows NAF to modulate its exploration–exploitation behavior over time. The output of the neural network is a categorical distribution over the discretized candidates, where each probability corresponds to the predicted acquisition value of selecting a specific $x_{s}$. During training, the next design candidate is sampled from this distribution, and the selected candidate is evaluated in the next iteration. During deployment, NAF directly selects the design candidate that has the highest acquisition value (value generated from NAF). We refer the readers to the original paper for more details about NAF (Volpp et al., 2019).
4. Summary and Limitations of NAF
NAF is a deep-learning-based meta-BO method with a key advantage: it can be trained with unlimited data without incurring computational overhead at deployment. Unlike GP-based approaches whose complexity grows with each new observation, NAF remains efficient regardless of dataset size. This makes NAF particularly well-suited for our HOMI, where synthetic users can be generated in virtually unlimited numbers via model sampling. Through such large-scale training, NAF learns acquisition policies that generalize across a wide range of synthetic user characteristics.
However, NAF has critical limitations when applied to real-world, human-involved optimization tasks. First, NAF is inherently designed to optimize for a single, fixed objective. In practice, many interactive systems involve multi-objective trade-offs, such as balancing speed and accuracy in input tasks, or comfort and visibility in interface layout. NAF lacks the flexibility to adjust its behavior when design objectives or user priorities shift. Second, NAF, as a deep learning model, performs well only when test-time tasks fall within the distribution of training tasks. NAF is likely to fail if real-world user behavior deviates significantly from the synthetic training distribution (i.e., constitutes an out-of-distribution task), which is likely given that human behaviors are inherently noisy and diverse.
Figure 2. Our Neural Acquisition Function+ (NAF+) has four main components. Here, we illustrate how they work together using keyboard adaptation as an example case. The surrogate model (Gaussian Process regression) captures the properties of the target user. The information of GP will then allow the neural acquisition function (pre-trained by synthetic data) and the typical acquisition function to generate a set of acquisition values. A novelty detector estimates how likely the new user is different from our training dataset based on the current observations; this information is then used to condition the aggregation of two acquisition functions. Finally, NAF+ suggests a design that is most likely to yield optimal user performance.
4.2. Overview of NAF+
To address the aforementioned challenges, we introduce NAF+, a new approach that builds upon NAF but incorporates two critical enhancements: the ability to dynamically adjust to multi-objective formulations, and fallback mechanisms for handling out-of-distribution users. These enhancements make NAF+ more flexible, generalizable, and better-suited for real-world human-in-the-loop applications. We assume that the target application involves multiple design objectives, which are eventually aggregated into a single scalar objective via a weighted-sum formulation; this is an approach commonly found in HCI work (Liao et al., 2024a; Bailly et al., 2013). Our contribution lies in training NAF+ such that it can support dynamic setting of objective weights during deployment without requiring any retraining, which allows the optimizer to respond to varying user needs or preferences on the fly. We also assume that while novel users (whose behavior significantly deviates from synthetic users) may appear during deployment, they can still benefit from standard BO mechanisms (e.g., exploration via typical acquisition functions like Expected Improvement). Therefore, NAF+ integrates a fallback mechanism to maintain robustness when facing novel, unseen users.
NAF+ has four key components to jointly achieve these enhancements: (1) GP-based surrogate model (GP): This component models the objective function based on the user’s observed performance data. It provides the predicted mean and variance necessary for computing acquisition values and remains the foundation for all downstream decisions. (2) Neural acquisition function (naf): This deep neural network replaces hand-crafted acquisition functions by learning a task-specific policy. Unlike the original NAF, our model is explicitly trained to condition on objective weights, enabling it to adjust its search strategy according to the prioritized trade-offs during deployment. To distinguish it from the original method, we refer to our version in lowercase: naf. Two more elements are introduced to mitigate the novel user challenge: (3) Expected Improvement (EI): A classical, well-established acquisition function used as a fallback. When the neural model performs poorly (e.g., due to encountering a novel user), EI provides a stable and exploration-oriented alternative. The acquisition values produced by naf and EI are combined via a weighted sum. (4) Novelty detector: This module estimates how novel the current user is with respect to the training distribution. Based on this assessment, it dynamically adjusts the weight between the neural acquisition score (naf) and the EI score. As novelty increases, more weight is placed on EI to ensure robust performance. In the below subsections, we will expand the details of each components.
4.3. Element 1: GP-based Surrogate model (GP)
The first component is the surrogate model of the target objective, reflecting the ongoing user. Following the standard BO framework, we employed a GP model to fit all the observed designs ($x$) and resulting objective function values ($y$). The information of GP (specifically, the predicted mean and variance at any given $x$) can be leveraged to compute acquisition values.
4.4. Element 2: Neural Acquisition Function (naf)
The naf model in NAF+ is a neural network trained to generate acquisition values for a specific interaction scenario, learned through interactions with synthetic users. We denote the naf model as $\alpha_{t,\theta}$, where $t$ indicates the current timestep and $\theta$ corresponsd to the neural network parameters. Our training process follows the general framework introduced in the original NAF paper (Volpp et al., 2019), with key extensions tailored to the requirements of HCI tasks.
First, we discretize the design space $\mathcal{X}$ into a fixed grid of candidates $\{x_{s}\}$. For each candidate $x_{i}$, where $i$ indexes the grid, the surrogate model (a Gaussian Process) provides a predicted mean $\mu_{t}(x_{i})$ and standard deviation $\sigma_{t}(x_{i})$ at optimization iteration $t$. These values form the core input features to naf, representing the model’s current belief about the design space. To enable dynamic adaptation over time, we include a budget-awareness signal: the proportion of the optimization process completed, calculated as $t/T$, where $T$ is the total optimization budget. This allows naf to shift its strategy from exploration in early iterations to exploitation in later stages.
A key innovation in NAF+ is its ability to support dynamic multi-objective optimization during deployment. To enable this, we append $N$ additional input values to the input vector, where $N$ is the number of design objectives. These values represent the weights assigned to each objective under the current trade-off configuration (e.g., $w_{1},w_{2},\dots,w_{N}$). For example, in a two-objective task with equal weights $[0.5,0.5]$, the input vector to naf will include these two values at the end, allowing the model to adapt its acquisition strategy accordingly. The full input to naf at iteration $t$ is thus:
(1) $\mu_{t}(x_{1}),\sigma_{t}(x_{1}),\dots,\mu_{t}(x_{s}),\sigma_{t}(x_{s}),t/T,w_{1},w_{2},\dots,w_{N}].$
The output of naf is a categorical distribution over the $s$ candidate points, denoted as $\text{CAT}[\text{nv}_{x_{1}},\text{nv}_{x_{2}},\dots,\text{nv}_{x_{s}}]$. Essentially, each $\text{nv}_{x_{i}}$ value represents the naf-generated acquisition value at a given design candidate $x_{i}$ across the full grid $x_{s}$. During training, a candidate design is sampled from this distribution to encourage exploration. During deployment, however, the design with the highest predicted acquisition probability is selected deterministically to maximize performance.
4.5. Element 3: Expected Improvement (EI)
Expected Improvement (EI) serves as the fallback acquisition strategy within NAF+. EI is one of the most widely adopted acquisition functions in Bayesian Optimization due to its simplicity, closed-form expression, and robust performance in balancing exploration and exploitation. Formally, EI at a candidate point $x$ is defined as:
(2) $\text{EI}(x)=\mathbb{E}[\max(0,f(x)-f(x^{+}))],$
where $f(x^{+})$ denotes the best observed objective value so far.
In NAF+, we apply EI to each discretized candidate $x_{i}\in\{x_{s}\}$ in the design space, producing a parallel set of acquisition scores $\text{EI}(x_{i})$. These scores are then combined with the values generated by the neural acquisition function (naf) to form the final acquisition values used for decision-making. The weighting between naf and EI is dynamically adjusted based on a novelty-aware mechanism, described in the following section, ensuring robustness when encountering out-of-distribution users.
4.6. Element 4: Novelty Detector
The novelty detector in NAF+ is a probabilistic module that estimates how likely a new user is to be out-of-distribution (i.e., significantly different from all previously seen synthetic users). Its primary role is to dynamically determine the weighting between the neural acquisition function (naf) and the fallback strategy (Expected Improvement, EI) when selecting the next design candidate.
At its core, the novelty detector is implemented as a Bayesian Neural Network (BNN), trained using the full set of $(x,y)$ pairs — design candidates and their corresponding objective function values — collected from synthetic users during pre-training. This training allows the BNN to approximate the population-level distribution of performance across diverse user characteristics, effectively modeling expected outcomes under a wide range of known conditions.
During deployment, when a new observation $(x,y)$ is collected from a real user, the BNN outputs a posterior predictive distribution for $x$, characterized by a mean $\mu_{t}(x)$ and standard deviation $\sigma_{t}(x)$. Using this, we compute the standardized z-score:
(3) $z=\frac{y-\mu_{t}(x)}{\sigma_{t}(x)}$
We then convert the z-score into a two-tailed p-value using the cumulative distribution function $\Phi(\cdot)$ of the standard normal distribution: $p=2\cdot\left(1-\Phi(|z|)\right)$. The p-value quantifies how surprising the real user’s performance is under the model trained on synthetic users. A low p-value (e.g., close to 0) indicates high novelty and suggests the user is likely out-of-distribution, whereas a high p-value indicates the user is well represented in the training data.
Over time, we accumulate multiple $(x,y)$ pairs from the same user and compute a running average of the p-values, denoted $\bar{p}$. We use this average to determine the contribution of EI in the acquisition strategy. Specifically, if $\bar{p}$ falls below a pre-defined novelty threshold $\tau$, we consider the user to be novel. The final acquisition score $A(x)$ used to select the next design candidate is computed as a weighted combination of the naf and EI scores:
(4) $A(x)=\lambda_{\text{EI}}\cdot\text{EI}(x)+(1-\lambda_{\text{EI}})\cdot\texttt{naf}(x)$
where the EI weight $\lambda_{\text{EI}}$ is computed as:
(5) $\lambda_{\text{EI}}=\begin{cases}1&\text{if }\bar{p}=0\\ 0&\text{if }\bar{p}\geq\tau\\ \frac{\tau-\bar{p}}{\tau}&\text{otherwise}\end{cases}$
This weighting mechanism ensures a smooth transition: when the user appears entirely novel ($\bar{p}=0$), full weight is given to EI; when the user aligns with the training distribution ($\bar{p}\geq\tau$), EI is ignored; otherwise, the influence of EI scales linearly based on the level of novelty.
4.7. Training and Deployment
Note that the EI and novelty detector components do not operate during the training phase. Only naf is trained by learning a policy to select promising design candidates based on inputs from the GP surrogate model. As discussed earlier, the training process is formulated as a RL problem, where the agent (i.e., naf) learns through repeated optimization episodes with synthetic users. Table 1 summarizes the RL formulation used in our meta-BO training pipeline. After training across diverse synthetic users, we obtain a large collection of design-performance pairs $(x,y)$ representing population-level behavior. These are then used to train the novelty detector (a Bayesian Neural Network, or BNN), which learns to estimate how likely a real user is out-of-distribution based on deviations from the synthetic population.
During deployment, all four components of NAF+ operate in concert. The surrogate model (GP) continuously fits the incoming data from real users. The naf model uses the updated GP predictions and the current objective weights to generate acquisition scores tailored to the user and design objectives. Simultaneously, the novelty detector assesses the novelty of the current user and dynamically adjusts the weight assigned to EI, which serves as a fallback acquisition function. The final acquisition value for each design candidate is computed as a weighted combination of naf and EI outputs, with the novelty-based weight $\lambda_{\text{EI}}$ governing the balance. We set a minimum iteration threshold $K$, so that the weighted-sum mechanism for acquisition values is activated only after $K$ iterations, ensuring that naf maintains full control during the initial phase.
RL Component NAF+ Equivalent
Policy $\pi_{\theta}$ Neural acquisition function (naf) $\alpha_{t,\theta}$
Episode Optimization episode with a synthetic user
Episode length $T$ Optimization budget $T$
State $s_{t}$ $\mu_{t}(x_{s}),\sigma_{t}(x_{s}),t/T,w_{1},\dots,w_{N}]$
Action $a_{t}$ Selecting a design candidate $x_{t}\in\mathcal{X}$
Reward $r_{t}$ $(f_{\text{max}}-f(x_{t}))$ (negative regret)
Transition $p(s_{t+1}|s_{t},a_{t})$ Noisy observation of $f(x_{t})$, GP update
Table 1. The RL formulation and mapping of NAF+.
4.8. Synthetic tests
Before conducting the user study, we designed two synthetic evaluations to systematically validate NAF+. These tests serve as controlled environments to probe capabilities that are difficult to isolate in live studies. Specifically, we focused on three goals: (G1) benchmarking against established baselines (standard BO and TAF), (G2) verifying that NAF+ can adapt to dynamic trade-offs between multiple objectives via explicit weight inputs, and (G3) evaluating robustness when encountering users whose behavior falls outside the training distribution. To cover both abstraction and realism, we used two tasks: a controlled double-Sphere benchmark (Test 1) for clean comparisons, and a semi-realistic typing simulation (Test 2) grounded in Fitts’ law and touch error models on touchscreen to mirror our target VR keyboard task in user study.
Key findings.
Across both tests, NAF+ consistently demonstrated advantages over baselines. For G1 (baseline comparison), it converged faster than standard BO and TAF, particularly in the early iterations where efficiency is most critical. For G2 (dynamic objective weighting), conditioning on explicit weight inputs enabled NAF+ to adapt its search strategy more effectively than the ablated variant without weights, confirming its ability to handle shifting priorities in multi-objective optimization. For G3 (robustness to novel users), the novelty-aware fallback proved essential: when synthetic users were sampled outside the training distribution, NAF+ recovered and improved steadily, while the no-fallback variant stalled and BO/TAF required more iterations to catch up. Together, these results establish that NAF+ combines sample efficiency, flexibility, and robustness, making it well-suited for real-world HILO where the weights on multiple objective can be adjusted dynamically and users can vary dramatically.
5. User Study: NAF+ for Mid-Air Keyboard Adaptation
We conducted a user study to evaluate the effectiveness of NAF+ in a classic HCI task: mid-air typing in virtual reality. Participants completed ten iterations of typing (each consisting of one to two sentences) with adaptive keyboards generated by three methods. Subsections 5.1–5.4 describe the study design, including the experimental conditions (methods), the user models used to train NAF+, the objective function, and the study apparatus. The study was conducted in two phases. Phase 1 (Subsection 5.5) provided data to pre-train two optimization methods, TAF and NAF+. Subsection 5.6 details the implementation and training procedure of NAF+. Phase 2, the final evaluation, is described in Subsections 5.7 (study setup) and 5.8 (results). Finally, Subsection 5.9 presents our discussion of the study findings.
5.1. Overall Study Design and Conditions
We compared three optimization methods: our proposed NAF+, Transfer Acquisition Function (TAF) (Wistuba et al., 2018; Liao et al., 2024a), and Continual Bayesian Optimization (ConBO) (Liao et al., 2025). Both TAF and ConBO represent state-of-the-art baselines. TAF adopts a meta-learning approach, where prior user data from complete HILO processes is leveraged to construct transferable models for new users. ConBO instead follows a continual learning approach: each user’s data directly serves as prior knowledge for subsequent users, without explicitly distinguishing between prior and target users.
To mitigate fatigue and learning effects, we constrained the study duration to one hour. Under this constraint, we limited the evaluation to three adaptation methods. Prior investigations have shown that standard Bayesian Optimization typically underperforms compared to approaches augmented with prior data (Liao et al., 2024b, 2025), so we excluded it as a baseline. Similarly, prior work demonstrated that manual calibration achieves performance comparable to prior-data–based optimization (Liao et al., 2025, 2024a), though in a substantially different setting where participants tuned the interface in advance. We therefore also excluded manual calibration from our study conditions.
Among the three selected methods, both NAF+ and TAF require pre-training, and the study was therefore organized into two phases. Phase 1 collected typing data to pre-train both methods, following protocols established in prior work (Bi and Zhai, 2016; Liao et al., 2024a). For NAF+, this data was used to fit the parameters of existing analytical typing models (Fitts’ Law and the DGD model) to our specific interaction context. While these models are well-established, no prior work has provided fitted parameters for the mid-air keyboard with direct touch; future work could bypass this step by reusing our fitted distributions. Also, other applications can potentially bypass this step and directly train NAF+ if reliable model parameters have already been estimated, such as touchscreen interactions where typing errors and movement time are well studied. For TAF, the same Phase 1 data was required to construct Gaussian Process (GP) models of prior users. Using a shared dataset for both methods ensures fairness: since TAF necessarily requires empirical data, we leverage the same data to calibrate NAF+.
5.2. Typing Simulation for Training NAF+
To pre-train NAF+, we implemented a generative simulation of sentence typing that models the motor execution and touch uncertainty of a human typist. The simulator combines two established models of touchscreen interaction: Fitts’ Law for movement time estimation (MacKenzie and Buxton, 1992) and the Dual Gaussian Distribution (DGD) model for touch accuracy (Bi et al., 2013; Bi and Zhai, 2016, 2013).
Each simulated user is parameterized by a 6-dimensional vector $\theta=[a,b,\alpha_{x},\sigma_{\alpha_{x}},\alpha_{y},\sigma_{\alpha_{y}}]$, which captures their motor and perceptual characteristics. The parameters $a$ and $b$ govern movement speed, while $\alpha$ and $\sigma$ terms control touch precision and variability in the horizontal and vertical directions.
Movement time.
The time to move between successive keystrokes is modeled using the classical Fitts’ Law equation:
(6) $MT=a+b\cdot\log_{2}\!\left(\frac{D}{W}+1\right),$
where $D$ is the Euclidean distance between the current and target key centers, and $W$ is the smaller dimension of the key.
Touch distribution.
Upon reaching a target key, the simulator samples a touch point from a 2D Gaussian distribution centered on the intended key:
(7) $\Sigma=\begin{bmatrix}\alpha_{x}\cdot W^{2}+\sigma_{\alpha_{x}}^{2}&0\\ 0&\alpha_{y}\cdot H^{2}+\sigma_{\alpha_{y}}^{2}\end{bmatrix}$.
This covariance formulation reflects the DGD model, where touch precision is influenced by both absolute noise ($\sigma$) and a size-dependent term ($\alpha$). Unlike prior work that analytically integrates over the distribution (Bi and Zhai, 2016), we simulate keystrokes via Monte Carlo rollouts: each touch sample is checked against the target key boundaries, and marked as correct or erroneous.
Sentence typing.
Test sentences are drawn from the Enron Mobile Email Dataset [^1] https://www.keithv.com/software/enronmobile/, with special characters removed and concatenated to a length at least 26 characters. Typing proceeds sequentially: the simulator begins at the spacebar, moves to each target key, samples a noisy landing point, and records both movement time and correctness. Additional Gaussian noise ($\sigma=0.15$ s) is added to capture motor variability. This procedure yields average movement time and error rate for a given keyboard layout. These settings remain identical in both our simulation experiments (Appendix B) and the user study.
5.3. Objective Function and Task Setting
The simulator outputs two primary metrics: typing speed (words per minute, WPM) and error rate. To combine them into a single scalar measure, we normalize each into unitless functions of Speed and Accuracy.
Through a 3-person pilot test, we determined that typing speed typically ranged between 5 WPM (worst) and 22 WPM (best). We therefore linearly normalized this range such that 5 WPM maps to 0 and 22 WPM maps to 1. Error rate, computed via Levenshtein distance, was observed to range between 0% (best) and 30% (worst). We similarly normalized this range so that 0% error maps to 1 and 30% to 0. While a 30% error rate may appear higher than in prior work, this upper bound is necessary in our adaptation setting, where small keyboards can lead to frequent errors.
Finally, speed and accuracy are combined in a weighted objective function:
(8) $\text{Objective}=w_{\text{speed}}\cdot\text{Speed}\;+\;w_{\text{accuracy}}\cdot\text{Accuracy},$
where we set $w_{\text{speed}}=0.7$ and $w_{\text{accuracy}}=0.3$ for the two phases in the user study. Other studies may adjust the weights to reflect different priorities.
In short, this setup captures the fundamental trade-off in mid-air typing: larger keys improve accuracy but increase movement times, while smaller keys accelerate typing but risk higher error rates. We therefore restrict the design space to keyboard widths and heights between 20 mm and 40 mm, and optimize within this range to maximize the objective in Equation 8.
5.4. Study Apparatus and Implementation
The virtual keyboard and typing experiment were implemented in Unity and executed on a Windows 10 desktop with an NVIDIA GeForce RTX 3050 GPU. We built the keyboard on top of the MRTK keyboard framework [^2] https://github.com/microsoft/MixedRealityToolkit-Unity, which we customized to support adaptive resizing of key height and width. The study was conducted on a Meta Quest Pro headset. Target sentences were sampled from the Enron Mobile Email Dataset, with special characters removed and multiple sentences concatenated to ensure a fixed length of 26 characters (excluding the space key) per iteration.
5.5. Phase 1: Gathering Data for NAF+ and TAF
The goal of Phase 1 was to collect training data to acquire typing models for NAF+ and Gaussian Process (GP) models for TAF.
5.5.1. Procedure
Five participants (2 female, 3 male; age 27–32) were recruited for Phase 1. Participants were instructed to type the presented sentences as quickly and accurately as possible on the mid-air keyboard. Before the main session, each participant completed 10 practice rounds with randomly varying keyboard dimensions to familiarize themselves with the adaptive setting. During the experiment, the keyboard remained at a fixed position (30 cm in front of the participant and tilted at $20^{\circ}$ relative to the horizontal plane), while only its width and height adapted. Each participant then typed 20 sentences, making the session last approximately 30 minutes, including practice.
To acquire data suitable for both TAF and NAF+, we adopted a hybrid procedure. In the first five iterations, keyboard dimensions were selected from evenly spaced values across the design space ($[20,25,30,35,40]$ mm) for both width and height, presented in randomized order. These evenly spaced settings ensure sufficient coverage for fitting typing error models (subsection 5.2). For the remaining 15 iterations, we used Bayesian optimization: the weighted objective (Speed + Accuracy) from prior trials was fitted to a surrogate model, and the acquisition function suggested new candidate designs. The full set of 20 iterations thus provided both uniformly sampled data for model fitting and adaptively sampled data for training GP priors in TAF.
Within each iteration, 26 key presses were recorded. We omitted the first six presses as a warm-up, leaving 20 effective key presses per iteration. Participants were instructed not to delete or correct typos. In cases where a participant attempted correction (e.g., typing “a” instead of “s,” then pressing “s” again), only the first (incorrect) press was retained for analysis.
5.5.2. Results
From the collected data, we constructed five GP models to serve as prior models for TAF. In parallel, we applied simple linear regression to approximate parameter values of the typing models. For each participant, individual parameter values were estimated and then summarized across participants as means and standard deviations.
For Fitts’ Law, the parameters were $a=0.164\pm 0.0352$ s and $b=0.39\pm 0.171$ s. For the error models, we obtained $\alpha_{x}=0.0148\pm 0.0011$, $\sigma^{2}_{x}=15.52\pm 2.093$, $\alpha_{y}=0.0133\pm 0.0011$, and $\sigma^{2}_{y}=15.93\pm 1.46$. These values provide empirical ranges for simulated users and serve as priors for model-based optimization in Phase 2.
5.6. Implementing and Training NAF+
Before Phase 2 of the user study, we prepared NAF+ by generating synthetic users, implementing its components, and training the model.
5.6.1. Creating Synthetic Users
Based on the parameter distributions estimated in Phase 1, we generated synthetic users by sampling from normal distributions over the six parameters ($a$, $b$, $\alpha_{x}$, $\sigma_{\alpha_{x}}$, $\alpha_{y}$, $\sigma_{\alpha_{y}}$). Each sampled parameter set corresponds to a unique synthetic user. Given a keyboard configuration, a synthetic user simulates typing behavior and outputs Speed and Accuracy, which are combined via the objective function (Equation 8). This approach enables large-scale training of NAF+without additional human data collection.
5.6.2. Implementation of NAF+
NAF+ consists of three components: the naf policy network, a BNN-based novelty detector, and a GP model. The naf network is a six-layer fully connected neural network, with 256 units per hidden layer and Tanh activations. The novelty detector is implemented as a two-layer fully connected Bayesian neural network (BNN) with 256 units per layer, using dropout ($p=0.2$) to approximate weight uncertainty. At inference time, multiple stochastic forward passes are performed to estimate predictive mean and variance, which are converted into z-scores and novelty p-values. The GP model is implemented using the botorch library as a single-task Gaussian Process [^3] https://botorch.org/docs/models/.
5.6.3. Training of NAF+
The naf network is trained using Proximal Policy Optimization (PPO) (Schulman et al., 2017). The actor and critic share the same architecture, with learning rates of $0.0003$ and $0.001$, respectively. PPO is run for $K=10$ epochs with a clipping parameter of 0.15. Although the user study employs fixed weights ($w_{\text{speed}}=0.7$, $w_{\text{accuracy}}=0.3$), we train naf to generalize across different weightings by randomly sampling $(w_{\text{speed}},w_{\text{accuracy}})$ during training. Synthetic users are dynamically generated throughout training, and the process continues until 80,000 steps are reached.
For the novelty detector, we consider 11 weight combinations by setting $w_{\text{speed}}\in\{0,0.1,0.2,\ldots,1.0\}$ and $w_{\text{accuracy}}=1-w_{\text{speed}}$. For each weight setting, we sample 20 synthetic users and evaluate 100 keyboard layouts (grid search), yielding 2,000 performance profiles per weight condition. These datasets are then used to train BNN models for novelty detection, each trained for 2,000 epochs.
5.7. Phase 2: Evaluation of Different Methods
In Phase 2, we evaluated three optimization methods for human-in-the-loop mid-air keyboard adaptation: our proposed NAF+, Transfer Acquisition Function (TAF), and Continual Bayesian Optimization (ConBO).
5.7.1. Study Design and Procedure
The study employed a within-subject design with 12 participants (5 female, age 24–33). Each participant completed three blocks, one per optimization method, with order counterbalanced using a Latin square. In each block, participants typed 10 sentences (iterations), following a short practice session of 10 iterations. Between blocks, a 5-minute break was provided to reduce fatigue. The entire session lasted less than 60 minutes per participant. Finally, the participants were asked to fill in the raw NASA-TLX questionnaire.
5.7.2. Methods
NAF+.
We described the implementation of NAF+ in Section 5.6. The novelty detector automatically selected the correct objective weight combination ($w_{\text{speed}}=0.7$, $w_{\text{accuracy}}=0.3$).
Transfer Acquisition Function (TAF).
TAF relies on a set of prior models—in our case, five GP models trained on the Phase 1 data. During deployment, acquisition function values are aggregated across these prior models using adaptive weights. The weights assigned to prior users decay according to two hyperparameters ($\alpha_{1},\alpha_{2}$), so that the influence of the new user’s model increases over time, allowing the optimization to converge toward user-specific preferences. We followed the parameterization of prior work (Liao et al., 2024a), setting $\alpha_{1}=4$ and $\alpha_{2}=0.2$.
Continual Bayesian Optimization (ConBO)
ConBO relies on a Bayesian neural network (BNN) population model trained on all previously observed users. Unlike TAF, ConBO does not reuse Phase 1 data; instead, the 12 participants in Phase 2 sequentially contribute training data to the population model. For example, data from Participant 1 updates the model, which then serves as prior knowledge for Participant 2, and so forth. As in TAF, the final acquisition function balances the population model with the current user’s model, assigning increasing weight to the latter to ensure personalized optimization. We largely followed the implementation of the original ConBO paper (Liao et al., 2025), with the same hyperparameters ($r_{0}=6$, $d_{r}=2$). To align with TAF, we set the decay parameters to $\alpha_{1}=4$ and $\alpha_{2}=0.2$.
5.7.3. Results
To evaluate optimization performance across iterations, we computed the objective function achieved in each iteration and method for every participant. The final objective function values for each condition (optimization method) at each iteration are shown in Figure 3. Note that we show and analyze the running best objective function values (i.e., the best performance achieved up to this iteration), which is a common practice for analyzing iterative optimization tasks in both machine learning and HCI fields. The raw typing speed (Figure 4) and the raw error rate (Figure 5) at each iteration are also shown, which jointly contribute to the final objective function.
We conducted a two-way repeated-measures ANOVA with factors Iteration (10 levels) and Method (NAF+, TAF, ConBO). There is a strong main effect of Iteration, $F(9,99)=27.16$, $p<.001$, indicating that participants’ performance, overall, improved over time. There was no significant main effect of Method, $F(2,22)=1.25$, $p=.307$, nor a significant Iteration $\times$ Method interaction, $F(18,198)=1.27$, $p=.211$. This indicates there is no method that is constantly better than the other throughout all iterations, which is aligned with the observation where three methods all converge toward the same optimal performance in later iterations (e.g., Iteration 6 onward). To further investigate whether different methods yield differences at different iterations, especially early iterations, we conduct one-way repeated-measures ANOVAs on each iteration separately. Significant effects of Method were found at Iteration 2, $F(2,22)=3.74$, $p=.040$, and Iteration 3, $F(2,22)=8.02$, $p=.002$, but not at any other iteration (all $p>.0.05$). Post-hoc pairwise comparisons with Bonferroni correction showed that at Iteration 2, NAF+ outperformed TAF, $t(11)=3.49$, $p_{\text{bonf}}=.015$, with no significant differences involving ConBO. At Iteration 3, NAF+ outperformed both TAF, $t(11)=3.52$, $p_{\text{bonf}}=.014$, and ConBO, t(11)=3.41t(11)=-3.41, $p_{\text{bonf}}=.018$, while TAF and ConBO did not differ significantly. The results show that while participants improved steadily across iterations regardless of method and all methods ultimately converge to comparable performances, NAF+ provided a performance advantage over TAF in earlier iterations (at Iterations 2–3) and over ConBO (at Iteration 3 only).
Finally, we analyzed participants’ responses to the NASA-TLX questionnaire and found no significant differences across conditions. This result is consistent with prior work (Lee et al., 2020; Liao et al., 2024a), which similarly reported that participants could not easily perceive differences between adaptation methods, even when objective performance differences were present. The primary reason was that within each adaptation method, different keyboard designs were proposed, making subjective comparison across methods challenging.
Figure 3. The weighted-sum performance achieved by different conditions (optimization methods) at each iteration. This plot presents the running best performance (i.e., the best performance of each participant achieved up to this iteration), which is a common way of analyzing iterative optimization tasks. The * sign indicates significant differences ($p<0.05$). The thicker, black line in each box shows the mean value, and the thin, orange line indicates the median value.
Figure 4. The typing speed at each iteration, which is then contributed to the calculation of the final objective function as described in subsection 5.3. The thicker, black line in each box shows the mean value, and the thin, orange line indicates the median value.
Figure 5. The error rate at each iteration, which is then contributed to the calculation of the final objective function as described in subsection 5.3. The thicker, black line in each box shows the mean value, and the thin, orange line indicates the median value.
5.8. Findings and Discussion
In this study, we investigated three optimization methods for HILO for mid-air keyboard adaptation. Particularly, we aim to understand how different strategies for leveraging prior knowledge influence performance over successive iterations. The key findings are that all three methods led to performance improvements across iterations, yet the resulting performances diverged in the early iterations: NAF+ provided an advantage in the second and third iterations.
TAF’s performance is significantly poorer in early iterations, which can be attributed to the limited coverage of its prior models. The five priors collected in Phase 1 could not sufficiently capture the diversity of population performance, leading to poor prediction in unexplored regions of the design space. In principle, expanding the number of priors could improve coverage and even approach optimal results. However, there are two fundamental challenges. First, there is no clear principle for predicting how many prior users are sufficient, which makes it difficult to guarantee generalization. Second, scaling up the number of priors linearly increases computation during deployment, rendering this approach less practical in real-world scenarios. By contrast, NAF+ has the unique strength of not being limited to a finite set of prior users. Instead, it can generate unlimited synthetic users by sampling from learned parameter distributions. This synthetic population enables extensive offline training before deployment, resulting in more consistent early-iteration performance. In addition, the novelty detector allows NAF+ to flexibly adjust the weighting between the NAF policy and the EI values calculated from the current user’s model. This mechanism provides robustness in out-of-distribution cases: when priors are less informative, NAF+ behaves similarly to TAF by relying more heavily on the current user’s data, ensuring at least comparable performance.
ConBO showed promising overall performance but was less effective in the earliest iterations, primarily because the first few participants exhibited relatively poor performance during their initial trials. The fundamental challenge of ConBO is that it learns gradually across users. For the first few participants, the population model contains little information, so its search behavior resembles random exploration. Once enough participants are aggregated, however, the model improves steadily. This trend is evident when comparing the performance of early and later groups of participants: Group 1 (P1–3) = 0.37 (std = 0.12), Group 2 = 0.61 (std = 0.07), Group 3 = 0.59 (std = 0.19), and Group 4 = 0.63 (std = 0.04). These results clearly show that ConBO benefits from accumulating user data, but at the expense of initial participants who experience suboptimal adaptation. In contrast, NAF+ does not suffer from this early-stage drawback, as its model is fully trained prior to deployment using synthetic users. The group-wise performances further illustrate this stability: Group 1 = 0.54 (std = 0.10), Group 2 = 0.56 (std = 0.12), Group 3 = 0.66 (std = 0.25), and Group 4 = 0.65 (std = 0.03). Thus, NAF+ provides stronger early performance and avoids the cold-start problem observed in ConBO.
In conclusion, this study highlights the trade-offs between different strategies for leveraging prior knowledge in HILO. TAF is limited by finite priors and scalability concerns, ConBO improves gradually but suffers from early-user effectiveness, and NAF+ combines the advantages of prior-based learning with scalable synthetic-user generation to achieve stronger early adaptation. These findings suggest that simulation-driven approaches such as NAF+ offer a promising path toward scalable, generalizable, and efficient human-in-the-loop optimization for adaptive interfaces.
6. Discussion
Optimization plays a key role in HCI, particularly for adaptive user interfaces. Prior work tends to follow one of two distinct paradigms: On one hand, model-based approaches leverage established user models to compute optimal designs, but they ignore real-time observations, unable to truly personalize to individual users or dynamic contexts. On the other hand, human-in-the-loop optimization adapts directly to individuals through online evaluation, but often requires significant exploration, making it slow to converge. In this work, we propose a novel concept, utilizing the advantages of both: pre-training optimizers using established user models to accelerate adaptation in human-in-the-loop settings. We define a novel conceptual framework, HOMI, which unifies model-based and interactive optimization by using synthetic users to train an adaptive optimizer prior to deployment. We provide the core components and stages of this framework and instantiate it through a concrete method: NAF+. NAF+ is a new optimization method that integrates a deep-learning-based acquisition function, trained via reinforcement learning on synthetic users to support efficient real-time adaptation. To extend its applicability to more realistic HCI scenarios, we build upon the Neural Acquisition Function (NAF) and introduce two key innovations: (1) the ability to handle dynamic trade-offs between multiple design objectives via weighted inputs, and (2) a novelty-aware fallback mechanism that detects novel users and adapts accordingly. Together, these contributions enable NAF+ to achieve fast, personalized adaptation in realistic, multi-objective design tasks. Our synthetic tests and user study validate the efficacy of our approach. Despite these promising results, several open questions and limitations remain.
Rethinking personalization: From model-fitting to meta-adaptation
Traditional model-based personalization requires fitting a parameterized user model to each new individual, which often involves complex and error-prone user inference pipelines before any adaptation can occur (Beaumont et al., 2002). Although recent learning-based inference approaches (e.g., amortized inference (Moon et al., 2023, 2024)) have been proposed to mitigate this burden, model inference remains a challenging step to support real-time adaptation to the diverse user characteristics. In contrast, HOMI adopts a different philosophy: we do not seek to model any single user perfectly. Instead, we train the optimizer on a broad distribution of synthetic users, enabling it to learn adaptation strategies that generalize across populations. At deployment, no explicit user modeling or inference is needed; the system simply adapts based on observed performance, guided by the meta-learned policy. This shift—away from high-fidelity modeling toward population-level diversity and meta-adaptation offers a more scalable and robust path for real-time personalization.
Re-purposing user models
Traditionally, user models have been developed to characterize and understand human performance, but rarely to serve a generative role. As a result, there has been little incentive to create or carefully fit such models for many interaction modalities, which explains why reliable parameter estimates are often missing, even when suitable models are available. This creates a fundamental challenge for our workflow: HOMI works best when models and parameters exist which allows our workflow to skip the parameter fitting step, yet in most realistic cases today, parameter fitting is likely to be needed. Our HOMI and NAF+ provide a new purpose for user models: shifting them from passive tools for understanding or characterizing human behavior, to the enabler of active interface optimization and adaptation. We hope this re-positioning will encourage the community to not only develop user models but also to document and share parameter values across different interactions and modalities, ultimately enabling broader deployment of synthetic-user-driven optimization.
Extending to other applications
Our approach is built upon the broad framework of BO and RL, which have shown strong scalability and generalizability across applications. In future work we will apply the HOMI framework to other interactive systems where well-established performance or cognitive models exist. Promising examples include menu optimization, gaze-based interfaces, and gesture input—scenarios where task-specific models can be leveraged to generate meaningful synthetic users and enable pretraining. We can then further validate the flexibility of our approach and demonstrate its potential as a general solution for scalable, adaptive interface optimization.
Utilizing advanced models to generate synthetic users
HOMI is grounded in the use of parameterized user models (e.g., Fitts’ law) to simulate synthetic users and enable large-scale pretraining. While this provides a scalable and principled way to approximate diverse user behaviors, it requires well-established models for a given interaction task. Such models are limited, and moreover, they are usually not available for more complex interactions involving rich cognitive processes or sequential decision-making or motions. Recent advances in Computational Rationality and reinforcement learning–based modeling of cognitive and motor processes  (Oulasvirta et al., 2022; Murray-Smith et al., 2022) offer promising alternatives for simulating realistic human behavior in such scenarios. Additionally, generative models, such as variational autoencoders, diffusion models, or large language models, are increasingly used to model human-like behavior from data, and could be adapted to synthesize user interaction patterns in tasks where explicit models are lacking (Li et al., 2024; Park, 2025). Future work could incorporate these approaches to expand the scope of HOMI, or explore hybrid user simulation strategies that blend analytical models with data-driven user representations. These directions can significantly broaden the applicability of our framework across a wider spectrum of HCI domains.
Generalization beyond synthetic users
Although our novelty detector provides a fallback mechanism for handling out-of-distribution users, the overall performance of NAF+ still depends on how well the synthetic user distribution approximates real-world variability. Moreover, the sampling of synthetic users presents an inherent trade-off. Sampling a broad and diverse range of synthetic users allows NAF+ to learn generalizable strategies across many user profiles, but this may come at the cost of reduced performance for any specific user. In contrast, sampling from a narrow range enables the optimizer to acquire more specialized strategies tailored to users that are more aligned with the synthetic ones, but at the risk of having a higher chance to encounter users outside of the training distribution. Balancing this exploration–exploitation trade-off in user simulation is a key open question. Future directions may include adaptive sampling, curriculum-based simulation, or progressive refinement of the synthetic user population based on deployment feedback. Such approaches could further enhance the generalization capacity of NAF+ in real-world human-in-the-loop scenarios.
Development of other HOMI methods
While this paper presents NAF+ as one realization of the broader HOMI framework, the framework itself is general and can support many alternative instantiations. One especially promising direction is to incorporate online user data into the meta-learning loop; that is, incorporating continual HILO (Liao et al., 2025) into our workflow. Currently, the meta-optimizer is trained exclusively on synthetic users generated from established models. However, as real user interaction data accumulates over time, integrating this data into the training process could allow the optimizer to continuously improve and better adapt to emerging usage patterns and behaviors. Such lifelong learning mechanisms would allow HOMI-based systems to evolve beyond their initial modeling assumptions and progressively refine their understanding of user diversity. Future HOMI methods could also explore alternative optimization strategies, such as moving beyond weighted-sum aggregation in multi-objective optimization and explicitly model Pareto frontiers, allowing users or systems to select among optimal trade-offs post hoc (Johns et al., 2023). Further, future research could investigate alternative acquisition architectures, such as transformer-based models or hybrid schemes that combine learned and analytical acquisition components. These extensions would further enhance the flexibility and impact of HOMI across a broader range of real-world HCI scenarios.
7. Conclusion
We introduce HOMI, a novel framework for HILO that leverages model-informed pre-training to accelerate real-time interface adaptation. By synthesizing large populations of simulated users from established user models, HOMI enables the training of powerful optimizers prior to deployment, bridging the gap between traditional model-based design and purely interactive approaches. To demonstrate the framework, we proposed NAF+, a concrete method that integrates a deep reinforcement learning–based acquisition function, dynamic multi-objective adaptation, and a novelty-aware fallback mechanism. Together, these components enable efficient and robust optimization across diverse users and changing objectives. Our results suggest that this new concept and approach jointly offer a scalable and generalizable path for future personalization systems: one that moves beyond per-user modeling and instead learns population-level adaptation strategies that generalize in deployment. We believe this work opens up a new direction for optimization in HCI, one that treats user models not just as analytical tools, but as generative engines for training adaptive, intelligent design systems. Finally, we hope this work provides a new lens for bridging in vitro (in simulation) with in situ (with real users) optimization, enabling more streamlined and adaptive future interfaces.
Appendix A Appendix: Overview of Simulated Tests
In this appendix, we describe the synthetic experiments conducted to validate the effectiveness of NAF+. These experiments are designed to address the following three goals:
  • Goal 1: Compare performance against baselines. We benchmark NAF+ against two established HCI optimization approaches: standard Bayesian Optimization (BO) and Transfer Acquisition Function (TAF).
  • Goal 2: Validate dynamic objective weighting. A key feature of NAF+ is its ability to adapt to varying weights across multiple design objectives during deployment. We test this capability under shifting weight settings.
  • Goal 3: Evaluate robustness to novel users. Another contribution of NAF+ is its mechanism for handling out-of-distribution users through a novelty-aware fallback strategy. We examine its efficacy in these scenarios.
To evaluate these goals, we conduct two synthetic tests. The first is a standard black-box optimization problem (Sphere function), commonly used for evaluating BO methods. The second is a semi-realistic simulation designed to mirror our target use case: optimizing mid-air keyboards for users with diverse performance profiles. In this simulation, we generate synthetic users with varying typing behaviors and aim to optimize keyboard layouts per individual.
A.1. Optimization Approaches (Conditions)
We define a set of optimization approaches, or “conditions,” which are held consistent across both synthetic tests.
1. NAF+(full).
This is the complete version of our proposed method, featuring both: (1) dynamic handling of multi-objective weights through explicit input, and (2) fallback support for novel users via the combination of EI and a novelty detector.
2. NAF+ without explicit multi-objective input.
This variant of NAF+ is trained to handle different objective weightings but does not receive explicit weight vectors as input during training or deployment. This tests the importance of conditioning the policy directly on objective weights.
3. NAF+ without novel user fallback.
This version excludes the components responsible for handling novel users (i.e., no EI fallback and novelty detection). It relies solely on the neural acquisition function (naf) trained on synthetic users.
4. Standard BO
This is the typical Bayesian Optimization setup, using Gaussian Processes and Expected Improvement as the acquisition function. It does not utilize any pretraining or meta-learning. This condition serves as a standard baseline in HCI for human-in-the-loop optimization.
5. Transfer Acquisition Function (TAF)
TAF is a meta-learning approach widely used in prior HCI work (Wistuba et al., 2018; Liao et al., 2024a). It builds a library of prior user models, each is as a standalone Gaussian Process (GP) trained on that user’s optimization data. During deployment, all prior models independently generate acquisition values for the candidate designs, and these values are aggregated—typically via a weighted sum—to guide the optimization for a new user. A key distinction between TAF and all variants of NAF+ lies in scalability. While TAF relies on a fixed set of prior users and its computational cost increases linearly with the size of the prior model library, NAF+ is trained on an arbitrarily large population of synthetic users but retains constant complexity. This enables NAF+ to scale more efficiently with increasing amounts of data.
Note that we did not include the vanilla NAF as a separate condition because it essentially operates as our NAF+ without both the explicit multi-objective input and the novel user fallback mechanism. The differences between NAF+ and NAF are already captured in the comparisons between Condition 1 vs Conditions 2 and 3.
A.2. Generating a Group of Synthetic Users
We follow standard procedures for evaluating meta-learning and transfer optimization methods (Volpp et al., 2019; Wistuba et al., 2018; Liao et al., 2024a), simulating a population of synthetic users who share core behavioral structure but exhibit individual variability. Each synthetic user is represented as a unique function, derived by modifying a shared base function through controlled transformations. We refer to these personalized variants as user functions. These are divided into a training group (used to train the meta-optimizer) and a test group (used to evaluate generalization at deployment).
For the first test, we determine a base function using a common test function for optimization tasks. Then, each user function is created by shifting the function (translating the input values before passing them to the function). The magnitude of the shifts is uniformly sampled from a distribution within a specified range, adding diversity among the users. Mathematically, the shift is represented as $x_{n}^{\prime}=x_{n}+\delta_{n}$, where $x_{n}$ is the original input, $x_{n}^{\prime}$ is the shifted input, $n\in{[1,N]}$ represents the parameters, and the shift amount $\delta_{n}\sim U(-\frac{shift\_range}{2},\frac{shift\_range}{2})$. This step simulates different user responses to the same design. We also scale the output of the function by a scalar factor, which is also sampled from a uniform distribution. This factor, denoted as $\mathcal{S}\sim U(1-\frac{scale\_range}{2},1+\frac{scale\_range}{2})$, introduces further diversity in users’ performance level. Through the combinations of these randomly sampled shift and scale, we generate a diverse set of user functions that preserve the structure of the base function while simulating realistic inter-user differences.
For the second test, we employ a user model based on existing literature, which has 6 user parameters. We sampled the parameter values from the range provided by the existing literature; a unique parameter assignment represents a specific user with a specific user performance given a design. More details will be provided in the second synthetic test (Appendix C).
A.3. General Procedures in the Synthetic Tests
Our synthetic tests follow a consistent structure consisting of three stages: training, testing, and evaluation on novel users.
Step 1: Training
We first train all NAF+ variations and TAF using a collection of prior user functions. For TAF, the training process involves fully optimizing each prior user function to build a set of GP models—one per user. To enable TAF to handle dynamic objective weights, we follow a prior approach (Liao et al., 2024a) where multi-objective optimization is performed during training. This results in GP models that can predict multiple objective values, allowing the optimizer to accommodate dynamic weight assignments during deployment. We train TAF with 15 prior user functions, which is consistent with earlier studies. Increasing the number of users further leads to prohibitively long computation times at deployment, as TAF’s runtime scales linearly with the number of stored models.
For all NAF+ variants, we dynamically generate synthetic users and continue training until a fixed number of training steps is reached ($80000$ steps for both synthetic tests). The three NAF+ variants share the same training procedure, with the only difference being that Condition 2 (NAF+ without explicit multi-objective input) is trained on multiple objective weight configurations but does not receive the weight information as part of the model input. Additionally, all collected $(x,y)$ pairs during training are used to train the Bayesian Neural Network–based novelty detector. Notably, the novelty detector is trained with the objective weights as part of the input, enabling it to condition its uncertainty estimation accordingly. However, Condition 3 (NAF+ without novelty detection) does not include this module and skips this training step.
Step 2: Testing
Next, we sample 20 test functions, each representing a new synthetic user with distinct characteristics. For each test function, the objective weights are randomly sampled from a uniform distribution. These functions are also shifted and scaled as described earlier to ensure diversity. Each condition is then deployed independently on these test functions to evaluate its performance with a budget of 20 iterations.
Step 3: Testing on Novel Users
To evaluate the robustness of each condition under distributional shift, we construct a separate set of “novel” user functions. These are generated using parameter values that lie completely outside the range observed during training. All optimization conditions are then tested on these novel users to assess their ability to generalize to unseen behaviors.
Appendix B Synthetic Test 1: Double-Sphere Functions
In this synthetic test, we define a composite 2D objective function to evaluate the optimizer’s ability to handle dynamic objective trade-offs as well as coping with novel users. Specifically, two sub-tests are conducted to evaluate two aspects of NAF+. The first sub-test (denoted as test 1.1) evaluates the importance of weight information, and the second sub-test (test 1.2) validates the importance of incorporating the typical expected improvement acquisition funciton.
B.1. Function Details
The function has two input parameters and two output objectives, each modeled as a 2D Sphere function centered at different locations. Specifically, the first Sphere function is centered at $(0.4,0.4)$, while the second is centered at $(0.6,0.6)$. Each Sphere function is defined as $y=1-\sum_{i=1}^{2}(\dot{x}_{i}-X_{i})^{2}\times\gamma$, where $X_{i}$ is the center coordinate and $\gamma=8$ controls the sharpness of the peak.
The final objective value is computed as the weighted sum of these two Sphere functions, with weights dynamically assigned during deployment. This creates a setting where the global optimum shifts depending on the weight configuration: when the first objective is prioritized, the optimal region lies near $(0.4,0.4)$; when the second is prioritized, it shifts toward $(0.6,0.6)$. To simulate user variability, we directly apply random shifts to the input space $\mathbf{x}$ in both dimensions, as well as scaling to the final weighted-sum output of the two sub-functions. Specifically, each synthetic user is defined by a unique shift vector and scaling factor, applied consistently across both objective components. This formulation introduces variation in both user preference (by altering the location of optima) and sensitivity (by modulating the steepness of the reward surface), while preserving the underlying structure of the task.
B.2. Configurations of each Condition
Here, we detailed the configurations of different conditions.
1. NAF+
We set the naf as a six-layer fully connected neural network, with each hidden layer containing 256 nodes and Tanh as the activation function. The naf is trained using Proximal Policy Optimization (PPO) (Schulman et al., 2017), where the critic network shares the same architecture as the actor (i.e., naf). The actor and critic learning rates are set to 0.0003 and 0.001, respectively. PPO is run for $K=10$ epochs with a clipping parameter of 0.15. To encourage generalization across varying objective weightings, the objective weights are randomly sampled during training. To be more specific, the weight on the first Sphere function, $\alpha$, is drawn uniformly from $[0,1]$, and the weight on the second is set to $1-\alpha$. This two-dimensional weight vector $(\alpha,1-\alpha)$ is also included as part of the input to the naf, allowing the acquisition function to adapt its strategy based on the current configuration. Note, when deploying on the test functions, the weights are given to NAF+ as a part of the input. We denote this method as NAF+ in the rest of this test.
2. NAF+ without weights information (NAF-w/o-Weight)
This condition is very similar to the full version of NAF+. The model setting, hyperparameter used in training, and training procedure are identical to NAF+. The only difference is that this method is not informed by the information of the objective weights.
3. NAF+ without EI (NAF-w/o-EI)
This condition is very similar to the full version of NAF+. The only difference is that there is no additional GPs constructed for the new function, and thus, there is no EI or novelty detector. The process is fully dominated by NAF all the time.
4. Transfer Acquisition Function (TAF)
This condition is Transfer Acquisition Function (denoted as TAF). This approach is a weighted-sum method; a series of GP models will be created and stored prior to deployment. We follow the previous works ((Liao et al., 2024a)) to configure TAF; 15 prior GPs are gathered, each is trained from a randomly sampled double-Sphere function, with the objective weights also randomly sampled. These GPs are constructed via standard BO, with 10 initial random sampling with 15 optimization steps. Similar to Liao et al. (2024a), we configure TAF with decaying model weights on the previous GPs, which allows the new GP model to gradually gain control over the optimization process. Specifically, $\alpha_{1}$ is set as 4 and $\alpha_{2}$ is set as 0.2.
5. Standard Bayesian optimization (BO)
The final condition is standard Bayesian optimization (BO). We used Expected Improvement as the acquisition function and a typical Single-Task GP implemented in BoTorch library as the surrogate model [^4] https://botorch.org/. We set BO with 6 random explorations followed by 14 optimization steps.
B.3. Result of Test 1.1 (Evaluating Weight Information)
Test 1.1 is designed to specifically evaluate the value of weight information. To this end, we sampled 20 test functions from the exact same range as the one used for training. Since the test functions are fully overlap with the training range, we exclude the condition with novel detection and additional EI (NAF-w/o-EI) from this test The log regret results are presented in Figure 6 left panel. Overall, we can observe that NAF+ started from an overall better performance and constantly outperformed other meta-BO approaches, such as NAF-w/o-Weight and TAF. Also, while TAF showed promising performance (most regret is less than $10^{-1}$), it still performed worse than NAF+ and NAF-w/o-Weight. The primary reason is that NAF+ is trained by nearly unlimited number of synthetic agents, allowing it to fully generalize across new tasks. On the other hand, TAF is based on a limited set of prior synthetic users. Lastly, within the first 10 iterations, all meta-BO approaches demonstrated better performance than standard BO, which starts from scratch. This highlights the benefit of prior experience.
Figure 6. The results of Test 1.1 (left) and Test 1.2 (right). Note that the value plot here is the $log_{10}$ of the regret.
B.4. Result of Test 1.2 (Evaluating Expected Improvement)
Test 1.2 evaluates the benefits of incorporating EI and a novelty detector to handle out-of-distribution cases. This time, we sampled 20 test functions outside the training range, which is a possible case in a real-world scenario where real users may not be fully aligned with our models. To that end, we directly shifted the center of the two Sphere functions from $(0.4,0.4)$ and $(0.6,0.6)$ to $(0.3,0.3)$ and $(0.7,0.7)$. This ensures the sampled functions will be different from the training functions, but not too distant. We include the condition with novel detection and additional EI (NAF-w/o-EI) from this test, but exclude the condition of weight information (NAF-w/o-Weight) since it is not a part of the main comparison. The log regret results are presented in Figure 6, right. Without incorporating EI based on the newly constructed GP, the performance of NAF did not make significant improvements throughout the iterations. This is expected, as the test functions are out of the training distribution. On the other hand, NAF+ did not start with a good performance due to the out-of-distribution test case. However, its performance gradually catches up with the other baselines, demonstrating the necessity of incorporating EI in real-world scenarios.
B.5. Summary of test 1
Test 1.1 answered goal 1 (comparing performance against baselines): we observe NAF+ and NAF-w/o-Weight generally outperformed TAF owing to its rich training data. NAF+ and NAF-w/o-Weight also significantly outperformed Standard BO in the first few iterations. Test 1.1 also answered goal 2 (validating dynamic objective weight): we found that overall, NAF+ converges faster than NAF-w/o-Weight, owing to its capability to take objective weight as input and adjust its search policy. Test 1.2 responded to goal 3 (evaluating robustness to novel users): we showed that when facing out-of-distribution users, the novelty detection with additional EI ensures robust convergence. Meanwhile, NAF-w/o-EI can not effectively address such scenarios.
Appendix C Synthetic test 2: Simulating softkeyboard typing
In the second synthetic test, we develop a typing user model that closely mirrors the target behavior of our real user study. This model predicts user performance on a given keyboard design, enabling us to evaluate design quality by simulating the act of typing a sentence on the sampled keyboard. As in the previous test, we conduct two sub-tests: one to assess the importance of including weight information as part of the model input, and another to evaluate the benefit of incorporating Expected Improvement (EI).
C.1. Model Details and Generating Synthetic Users
In this test, a keyboard design is defined by its spatial dimensions, including height and width. The user model predicts two key performance metrics: (1) typing error, defined as the probability of pressing an incorrect key, and (2) movement time between consecutive key presses. The model is composed of two analytical components: Fitts’ Law (MacKenzie and Buxton, 1992) for modeling movement time, and the Dual Gaussian Distribution (DGD) model (Bi and Zhai, 2016) for predicting touch error on touchscreen keyboards. Further implementation details are provided in Section 5 of the main paper.
To generate synthetic users, we assign values to the model parameters $a,b,\alpha_{x},\sigma_{\alpha_{x}},\alpha_{y},\sigma_{\alpha_{y}}]$ by sampling from distributions informed by prior work. Bi et al. (2013) report Fitts’ Law parameters fitted to touchscreen typing as $a=144.3$ and $b=75.636$, with the unit being milliseconds. We use these values as the means of two independent Gaussian distributions (each with a standard deviation of 15) and sample from them to generate user-specific values of $a$ and $b$. Here, we emphasize that this is a synthetic task; the exact parameter values vary for different input modalities and keyboard settings, and here we only take the available model values to demonstrate the efficacy of our approach.
For the touch error model, we adopt the generic parameterization from the DGD model (Bi and Zhai, 2016), setting the means of the Gaussian distributions to $\alpha_{x}=0.0075$, $\sigma_{\alpha_{x}}=1.296$, $\alpha_{y}=0.0108$, and $\sigma_{\alpha_{y}}=1.153$. The corresponding standard deviations are arbitrarily set to $0.001$, $0.01$, $0.001$, and $0.01$, respectively. Each parameter is sampled independently, resulting in a diverse population of synthetic users with varying motor behavior and touch precision profiles.
C.2. Configurations of each Condition
All the conditions remain the same as in the previous synthetic test (subsection B.2).
C.3. Result of Test 2.1 (Evaluating Weight Information)
Similar to the previous test, Test 2.1 is designed to evaluate the value of the weight information. We sampled 20 test functions from the exact same range as the one used for training. We again exclude the condition with novel detection (NAF-w/o-Weight) from this test as there will not be novel functions. The log regret results are presented in Figure 7, left panel. NAF+ and NAF-w/o-Weight start from an overall better performance than the others. Especially, NAF+ converge to optimal performance around the second iteration, and NAF-w/o-Weight converges around the fifth iteration, showing NAF+ still benefits from the additional weight information as a part of the input. We observed that TAF with 15 prior models struggled to improve significantly in this case. This might be because of the higher number of parameters (2 in Test 1 and 6 in Test 2). To cover this number of dimensions, potentially more models will be required. Finally, same as the previous test, Standard BO starts from a worse performance and gradually converges toward optimal performance.
C.4. Result of Test 2.2 (Evaluating Expected Improvement)
Test 2.2 evaluates the value of merging EI and the novelty detector. To create a scenario where the new user is completely outside of the training range, we significantly shift the range of sampling the model parameters. The mean for sampling synthetic users are set as $[180,120,0.0145,1.6,0.03,2.0]$ for $a,b,\alpha_{x},\sigma_{\alpha_{x}},\alpha_{y},\sigma_{\alpha_{y}}]$, respectively, and the standard deviation remains the same as. These ranges are significantly different from the parameter settings for training, ensuring all sampled users are novel. We sampled 20 test functions from the test range. The log regret results are presented in Figure 7, right panel. We find that NAF+ and NAF-w/o-EI start from overall better performance, but NAF-w/o-EI does not improve significantly throughout the iterations, while NAF+ continually improves. This, again, highlights the potential issue of not incorporating an additional EI to handle out-of-distribution cases: naf can not generalize to other cases once it is fully trained to a certain range. TAF and standard BO perform well in this task. However, they start from a worse performance. BO, especially, requires 4 to 5 iterations to catch up with other conditions.
C.5. Summary of Test 2
Test 2.1 addresses goal 1 (comparing performance against baselines). We observe that NAF+ and NAF-w/o-Weight consistently outperform TAF, likely due to the higher dimensionality of the input space in this test. TAF, relying on a limited set of prior models, would require a larger number of samples to adequately cover the expanded design space. NAF+ and NAF-w/o-Weight also significantly outperform standard BO in the early iterations, demonstrating the benefits of leveraging prior knowledge. Test 2.1 also supports goal 2 (validating dynamic objective weighting). Compared to NAF-w/o-Weight, NAF+ achieves faster convergence by conditioning its acquisition strategy on the provided objective weights, resulting in higher sample efficiency. Test 2.2 addresses goal 3 (evaluating robustness to novel users). We find that NAF-w/o-EI struggles to adapt when users fall outside the training distribution, while NAF+ —equipped with a novelty-aware EI fallback—successfully recovers and improves performance over time.
In summary, both Test 1 (a controlled synthetic benchmark) and Test 2 (a more realistic simulation grounded in HCI literature) converge on the same conclusion: NAF+, through its use of explicit weight input and novelty-aware EI, achieves greater sample efficiency and robustness across a wide range of user profiles. The positive outcomes of Test 2 further motivate our user study, which targets typing performance on mid-air keyboards, highly relevant with Test 2 – a touch-based keyboard for mobile devices.
Figure 7. The results of Test 2.1 (left) and Test 2.2 (right). Note that the value plot here is the $log_{10}$ of the regret.